JNJ, Inc., is a fast-growing technology company. Management projects rapid growth of 30 percent for the next two years, then a growth rate of 17 percent for the following two years. After that, a constant-growth rate of 8 percent is expected. The firm expects to pay its first dividend of $2.45 a year from now. If dividends will grow at the same rate as the firm and the required rate of return on stocks with similar risk is 22 percent, what is the current price of the stock?
Answer #1

Expected Dividend in 1 year, D1 = $2.45

Growth rate for next 2 years is 30%, for next two years is 17% and a constant growth rate (g) of 8% thereafter.

D2 = $2.45 * 1.30 = $3.185
D3 = $3.185 * 1.30 = $4.1405
D4 = $4.1405 * 1.17 = $4.8444
D5 = $4.8444 * 1.17 = $5.6679
D6 = $5.6679 * 1.08 = $6.1213

Required Return, rs = 22%

P5 = D6 / (rs - g)
P5 = $6.1213 / (0.22 - 0.08)
P5 = $43.7236

P0 = $2.45/1.22 + $3.185/1.22^2 + $4.1405/1.22^3 + $4.8444/1.22^4 + $5.6679/1.22^5 + $43.7236/1.22^5
P0 = $26.89

So, the current price of the stock is $26.89
Answer #2

D3=$3.185*1.17=$3.772645

starting in yr 3 dividend growth rate is 17%.